OpenStudy (anonymous):

solve these system by graphing yx-3

OpenStudy (anonymous):

Plot the lines y= x and y = x - 3 on graph paper. Then shade the unwanted regions ie the part showing y > x and the part showing y <x-3 The values for y in the unshaded region are your solutions This is also called linear programming

OpenStudy (anonymous):

\[ylexit is like this i couldn't type the less than or equal \to: \]

OpenStudy (anonymous):

\[y \le x\] y>x-3

OpenStudy (anonymous):

\[ x+y-3\le0\] \[y-2\le\]